# A 99-Line Homogenization Code for Lattice-skin Plate Structures

### 25 Apr 2026 v1

@@ 2.3 LPS-H framework for finite-thickness evaluation
In numerical homogenization, the $\mathbf{ABD}$ matrix is unknown, so a column-by-column probing strategy is adopted. According to the rule of matrix multiplication, if an orthogonal unit deformation basis vector $\mathbf{E}^{(I)}$ is applied such that only its $I$th component is equal to 1 and all other components are zero, then the resulting internal-force response vector $\mathbf{R}^{(I)}$ is numerically equal to the $I$th column of the $\mathbf{ABD}$ matrix. Therefore, solving the full $6\times 6$ matrix can be converted into solving six independent microscopic response problems under six macroscopic unit deformation modes, where $I=1,2,\dots,6$.
@@
After Eq. (7) is applied to the RVE, the initial macroscopic strain field disturbs the local equilibrium of the microscopic solid because the metamaterial contains an interwoven topology of solid matrix and voids. To re-establish equilibrium, the structure develops a local fluctuating displacement field, together with the fluctuating strain $\boldsymbol{\varepsilon}^{fluc(I)}(\mathbf{x})$. By solving the finite element weak form under the mixed boundary conditions of 2D-PBC and 1D-FBC, this fluctuating strain can be obtained, and the microscopic local in-plane stress field $\boldsymbol{\sigma}_{p}^{(I)}(\mathbf{x})$ for the $I$th loading case can be recovered from the local constitutive tensor $\mathbf{C}(\mathbf{x})$ of the solid material:
@@
As described in Step 1, the resulting internal-force vector corresponds numerically to the $I$th column of the $\mathbf{ABD}$ matrix. In this way, a mapping is established from microscopic stress integration to macroscopic stiffness components:
@@ 27 B[5, 0::3], B[5, 1::3] = dN[1], dN[0]
Lines 17-30 show the analytical calculation of the element stiffness matrix $\mathbf{K}_{e}$. To evaluate the stiffness matrix on a regular voxel grid, the program introduces an isoparametric mapping from the physical coordinate system $(x,y,z)$ to the dimensionless local natural coordinates $(\xi,\eta,\zeta\in[-1,1])$. For an orthogonal regular element with side lengths $dx$, dydy, and $dz$, the Jacobian matrix $\mathbf{J}$ degenerates into a diagonal matrix, and its determinant is constant over the entire element domain:
@@ 3.3 Loop-Free Assembly of Global Matrices
For Element $e$, under the $I$th deformation case, the equivalent nodal load induced by the initial macroscopic strain is written as
@@ 4.2 Mesh Convergence Analysis
In voxel-based finite element homogenization, the mesh resolution affects both the geometric fidelity of the implicit surface boundary and the number of DOFs in the discrete model. To balance computational accuracy and efficiency, a mesh convergence analysis is carried out for the LPS-H framework. The sheet-type Primitive structure is selected as the reference case, while the cell array and the relative density of 0.15 are kept unchanged. The one-dimensional mesh resolution of the unit cell, $N$, is increased from 20 to 125 with a sampling interval of 5. The in-plane principal tensile stiffness $A_{11}$ and the out-of-plane principal bending stiffness $D_{11}$ are extracted from the effective stiffness matrix at each resolution. Their convergence histories are shown in Figure 5.
@@
The results show that when the one-dimensional mesh resolution is low, the predicted in-plane and out-of-plane stiffness values exhibit noticeable fluctuations. This is related to the stair-step modeling effect produced when a coarse voxel grid approximates a complex TPMS surface. As the mesh resolution increases, the voxel model approximates the implicit surface more accurately and the stair-step effect gradually weakens. When $N$ exceeds 60, the values of $A_{11}$ and $D_{11}$ become stable and show asymptotic convergence.
@@ 5.1 Extension to Multimaterial Lattice-skin Plate Structures
More generally, the material distribution inside many multimaterial structures is irregular. In this case, an external material-label matrix material_grid, consistent with the structural geometry and of size $N_{x}\times N_{y}\times N_{z}$, is usually maintained. Here, the value $1$ represents the first material, $2$ represents the second material, and so on. In the code, Boolean masking can be used to extract the material IDs of the active elements, and np.select can then realize loop-free mapping of multiphase material properties.
@@ Appendix B Comparison of LVS-H, LPS-H, and the Full-Scale Model
This appendix presents the static comparison results for a representative Lattice-skin Plate Structure with dimensions of $120\times 120\times 10$ mm under three modeling strategies: the full-scale finite element model, LPS-H, and LVS-H, as shown in Figure 9. To ensure fair comparison, the three models use the same boundary conditions and loading setup: the upper boundary along the $y$ axis is fully fixed, a displacement load of $1$ mm is applied to the lower boundary, and the overall support reaction of the structure is used as the effective stiffness indicator.

### 01 May 2026 v2

@@ 2.3 LPS-H framework for finite-thickness evaluation
In numerical homogenization, the $\mathbf{ABD}$ matrix is unknown, so a column-by-column probing strategy is adopted. According to the rule of matrix multiplication, if an orthogonal unit deformation basis vector $\mathbf{E}^{(I)}$ is applied such that only its IIth component is equal to 1 and all other components are zero, then the resulting internal-force response vector $\mathbf{R}^{(I)}$ is numerically equal to the IIth column of the $\mathbf{ABD}$ matrix. Therefore, solving the full $6\times 6$ matrix can be converted into solving six independent microscopic response problems under six macroscopic unit deformation modes, where $I=1,2,\dots,6$.
@@
After Eq. (7) is applied to the RVE, the initial macroscopic strain field disturbs the local equilibrium of the microscopic solid because the metamaterial contains an interwoven topology of solid matrix and voids. To re-establish equilibrium, the structure develops a local fluctuating displacement field, together with the fluctuating strain $\boldsymbol{\varepsilon}^{fluc(I)}(\mathbf{x})$. By solving the finite element weak form under the mixed boundary conditions of 2D-PBC and 1D-FBC, this fluctuating strain can be obtained, and the microscopic local in-plane stress field $\boldsymbol{\sigma}_{p}^{(I)}(\mathbf{x})$ for the IIth loading case can be recovered from the local constitutive tensor $\mathbf{C}(\mathbf{x})$ of the solid material:
@@
As described in Step 1, the resulting internal-force vector corresponds numerically to the IIth column of the $\mathbf{ABD}$ matrix. In this way, a mapping is established from microscopic stress integration to macroscopic stiffness components:
@@ 27 B[5, 0::3], B[5, 1::3] = dN[1], dN[0]
Lines 17-30 show the analytical calculation of the element stiffness matrix $\mathbf{K}_{e}$. To evaluate the stiffness matrix on a regular voxel grid, the program introduces an isoparametric mapping from the physical coordinate system $(x,y,z)$ to the dimensionless local natural coordinates $(\xi,\eta,\zeta\in[-1,1])$. For an orthogonal regular element with side lengths dxdx, dydy, and dzdz, the Jacobian matrix $\mathbf{J}$ degenerates into a diagonal matrix, and its determinant is constant over the entire element domain:
@@ 3.3 Loop-Free Assembly of Global Matrices
For Element $e$, under the IIth deformation case, the equivalent nodal load induced by the initial macroscopic strain is written as
@@ 4.2 Mesh Convergence Analysis
In voxel-based finite element homogenization, the mesh resolution affects both the geometric fidelity of the implicit surface boundary and the number of DOFs in the discrete model. To balance computational accuracy and efficiency, a mesh convergence analysis is carried out for the LPS-H framework. The sheet-type Primitive structure is selected as the reference case, while the cell array and the relative density of 0.15 are kept unchanged. The one-dimensional mesh resolution of the unit cell, NN, is increased from 20 to 125 with a sampling interval of 5. The in-plane principal tensile stiffness $A_{11}$ and the out-of-plane principal bending stiffness $D_{11}$ are extracted from the effective stiffness matrix at each resolution. Their convergence histories are shown in Figure 5.
@@
The results show that when the one-dimensional mesh resolution is low, the predicted in-plane and out-of-plane stiffness values exhibit noticeable fluctuations. This is related to the stair-step modeling effect produced when a coarse voxel grid approximates a complex TPMS surface. As the mesh resolution increases, the voxel model approximates the implicit surface more accurately and the stair-step effect gradually weakens. When NN exceeds 60, the values of $A_{11}$ and $D_{11}$ become stable and show asymptotic convergence.
@@ 5.1 Extension to Multimaterial Lattice-skin Plate Structures
More generally, the material distribution inside many multimaterial structures is irregular. In this case, an external material-label matrix material_grid, consistent with the structural geometry and of size $N_{x}\times N_{y}\times N_{z}$, is usually maintained. Here, the value 11 represents the first material, 22 represents the second material, and so on. In the code, Boolean masking can be used to extract the material IDs of the active elements, and np.select can then realize loop-free mapping of multiphase material properties.
@@ Appendix B Comparison of LVS-H, LPS-H, and the Full-Scale Model
This appendix presents the static comparison results for a representative Lattice-skin Plate Structure with dimensions of $120\times 120\times 10$ $m$ under three modeling strategies: the full-scale finite element model, LPS-H, and LVS-H, as shown in Figure 9. To ensure fair comparison, the three models use the same boundary conditions and loading setup: the upper boundary along the $y$ axis is fully fixed, a displacement load of 11 mm is applied to the lower boundary, and the overall support reaction of the structure is used as the effective stiffness indicator.